# TRACE: A Framework for Analyzing and Enhancing Stepwise Reasoning in Vision-Language Models

### v2

@@ 5 Conclusion and Future Work
In the Cartesian coordinate system xOyxOy, a circle centered at the origin $O$ passes through the point $A(13,0)$. The line y=kx3k+4y=kx-3k+4 intersects the circle $\odot O$ at points $B$ and $C$. What is the minimum length of the chord BCBC?
@@
Main Question: What is the minimum length of chord BCBC? $6\sqrt{17}$ / 24
@@
In $\triangle ABC$, $E$ is the midpoint of ACAC, and $D$ lies on BCBC such that $BD:CD=2:3$. Lines ADAD and BEBE intersect at $O$. If $S_{\triangle AOE}-S_{\triangle BOD}=1$, find the area of $\triangle ABC$.
@@
Q1: What is the ratio of BDBD to CDCD? $(2/3)$ / $(2/3)$
Q2: Is $E$ the midpoint of ACAC? Yes / Yes
@@
The model first fails at Q3, incorrectly computing the ratio of areas $\triangle AOE:\triangle BOD$ as 44 instead of $5/4$. This miscalculation propagates to subsequent area computations, leading to an incorrect area for $\triangle ABC$.
Why Area Ratios Determine the Final Answer The ratio $\dfrac{S_{\triangle AOE}}{S_{\triangle BOD}}$ links the condition $S_{\triangle AOE}-S_{\triangle BOD}=1$ to the absolute area of $\triangle ABC$. Geometric Relationship: Intersection of cevians ADAD and BEBE defines proportional sub-triangles. Correct ratio ($5/4$) allows solving for $S_{\triangle AOE}$ and $S_{\triangle BOD}$. Summing sub-triangle areas scales to $S_{\triangle ABC}=10$. ARS Dependency Justification: Q3 directly governs the chain of reasoning. An incorrect ratio leads to errors in all subsequent area computations.
@@
In rectangle OABCOABC, OAOA and OCOC lie on the $x$- and $y$-axes. $B=(3,2)$, points $D$ and $E$ lie on ABAB and BCBC with $BD=BE=1$. Folding $\triangle BDE$ along DEDE maps $B$ to $B^{\prime}$. Find the coordinates of $B^{\prime}$.
@@
Q2: Length of BDBD? 11 / 1
Q3: Length of BEBE? 11 / 1
@@
Q6: Slope of DEDE? Undefined / 1-1
Q7: Equation of line DEDE? $x=3$ / y=x+4y=-x+4

### v3

@@ 5 Conclusion and Future Work
In the Cartesian coordinate system $xOy$, a circle centered at the origin $O$ passes through the point $A(13,0)$. The line y=kx3k+4y=kx-3k+4 intersects the circle $\odot O$ at points $B$ and $C$. What is the minimum length of the chord $BC$?
@@
Main Question: What is the minimum length of chord $BC$? $6\sqrt{17}$ / 24
@@
In $\triangle ABC$, $E$ is the midpoint of $AC$, and $D$ lies on $BC$ such that $BD:CD=2:3$. Lines $AD$ and $BE$ intersect at $O$. If $S_{\triangle AOE}-S_{\triangle BOD}=1$, find the area of $\triangle ABC$.
@@
Q1: What is the ratio of $BD$ to $CD$? $(2/3)$ / $(2/3)$
Q2: Is $E$ the midpoint of $AC$? Yes / Yes
@@
The model first fails at Q3, incorrectly computing the ratio of areas $\triangle AOE:\triangle BOD$ as $4$ instead of $5/4$. This miscalculation propagates to subsequent area computations, leading to an incorrect area for $\triangle ABC$.
Why Area Ratios Determine the Final Answer The ratio $\dfrac{S_{\triangle AOE}}{S_{\triangle BOD}}$ links the condition $S_{\triangle AOE}-S_{\triangle BOD}=1$ to the absolute area of $\triangle ABC$. Geometric Relationship: Intersection of cevians ADAD and $BE$ defines proportional sub-triangles. Correct ratio ($5/4$) allows solving for $S_{\triangle AOE}$ and $S_{\triangle BOD}$. Summing sub-triangle areas scales to $S_{\triangle ABC}=10$. ARS Dependency Justification: Q3 directly governs the chain of reasoning. An incorrect ratio leads to errors in all subsequent area computations.
@@
In rectangle $OABC$, OAOA and $OC$ lie on the $x$- and $y$-axes. $B=(3,2)$, points $D$ and $E$ lie on $AB$ and $BC$ with $BD=BE=1$. Folding $\triangle BDE$ along $DE$ maps $B$ to $B^{\prime}$. Find the coordinates of $B^{\prime}$.
@@
Q2: Length of $BD$? 11 / 1
Q3: Length of $BE$? 11 / 1
@@
Q6: Slope of $DE$? Undefined / 1-1
Q7: Equation of line $DE$? $x=3$ / y=x+4y=-x+4